The perimeter of a rectangle is 72 in. The base is 3 times the height. Find the area of the rectangle.
Answer by JulietG(1812)

The perimeter is a fence around. It consists of 2 pieces of length and 2 pieces of width.
2L + 2W = 72
W = 3L (we'll call the base the 'width', and the height the 'length')
Substitute the value of W in the second equation into the first.
2L + 2(3L) = 72
Multiply
2L + 6L = 72
Add
8L = 72
Divide each side by 8
L = 9
If the length is 9, then the width is 27 (3 times the height).
.
Let's check the answer.
For a perimeter fence with 2 pieces at 27 and 2 pieces at 9, we have
27+27+9+9 = 72
72 = 72
Success!
.
Now that we know the length and width, we can find the area by multiplying them.
9" * 27" = 243 square inches
(Check my math to be certain)